OpenStudy (anonymous):

A student club has 12 members. How many ways can the club choose a president, vice president, secretary, and treasurer?

Parth (parthkohli):

^ lol

OpenStudy (anonymous):

12x11x10x9

OpenStudy (anonymous):

hey thats no nice @wwsbwilson

Parth (parthkohli):

12c3

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {12! \over (12 - 3)! 3!} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {12 \cdot 11 \cdot 10 \cdot 9 \over 3 \cdot 2 \cdot 1} }\)

OpenStudy (anonymous):

11 880

Parth (parthkohli):

You can cancel, or use a calculator

OpenStudy (anonymous):

11,880

OpenStudy (anonymous):

The bowling team at Lincoln High School must choose a president, vice president, and secretary. If the team has 10 members, which expression could be used to determine the number of ways the officers could be chosen?

Parth (parthkohli):

Oh wait

Parth (parthkohli):

(12 * 11 * 10) / 6

Parth (parthkohli):

1320/6

OpenStudy (anonymous):

How many license plates can be made which have the form ##LLL (2 numbers followed by 3 letters with repetition allowed)?

OpenStudy (anonymous):

@mathslover can u help